# Privacy-Preserving In-Context Learning with Differentially Private Few-Shot Generation

### v1

@@ 1 Introduction
The emergence of in-context learning (ICL) with large language models (LLMs), popularized by the seminal work of Brown et al. [7], has revolutionized the field of natural language processing and machine learning; see [12] for a survey on ICL and the references therein. In-context learning involves downstream task adaptation without modifying a pre-trained model’s weights. This is achieved by conditioning the model through a series of demonstrations of the task at hand appended as a prompt. An advantage of ICL is that it offers a cost-effective and adaptable alternative to fine-tuning LLMs. By leveraging the model’s pre-trained knowledge, it enables efficient generalization across tasks, allows for quick adaptation to new domains or concepts, and requires only a handful of labeled examples for adaptation.
However, privacy is a concern when deploying LLMs with users’ data incorporated into prompts. As an example, consider healthcare AI applications, where clinical reports belonging to the patients may be used as demonstrations to provide relevant context to the LLM to answer queries. A malicious adversary might attempt to circumvent API restrictions through jailbreaking thereby gaining direct access to the demonstrations as depicted in Figure 1.111As also observed in a real-world system built on an LLM: see https://tinyurl.com/nhzadhuz. More generally, it is a major concern that LLMs may regurgitate prompt data in their output [33, 14, 38]. These scenarios raise privacy risks regarding the data used for constructing the prompt. To mitigate these risks associated with prompt data, heuristic precautions can be taken such as removing personal identifiable information (PII). Unfortunately, it is well known in the privacy literature that a document without any PII may still be linked to an individual with auxiliary information an adversary possesses, which will inevitably lead to privacy violations (e.g., based on GDPR [4]).
Figure 1: Description of a potential privacy violation when few-shot demonstrations are pulled from a private dataset in an ICL framework for a healthcare application. A malicious adversary attempts a basic prompt injection attack and gains direct access to the demonstrations. Basic heuristics such as personal identifiable information (PII) removal may still leave linkable information [4] to an individual in case the adversary has auxiliary information (e.g., a unique patient with a particular disease or treatment) and do not prevent against privacy violations.
In this context, differential privacy (DP) [17] is widely regarded as the gold standard in safeguarding the privacy of individuals contributing their data to various applications, ensuring that their sensitive information remains confidential while still enabling valuable insights to be derived from the aggregated data [36].
@@ 1.1 Our Contributions
  • We introduce a new algorithm to perform ICL with DP guarantees on private datasets (Figure 2). Our algorithm generates synthetic few-shot demonstrations from the original private dataset to be used for ICL during inference, and guarantees DP with respect to examples in the private dataset. At a high level, our approach utilizes the capabilities of LLMs in terms of generating a data sample similar to the ones in the original dataset based on demonstrations of a number of samples from the private data. To guarantee DP on the private dataset, we perform this operation by privately aggregating the generation process coming across disjoint subsets of the private data.
  • We empirically evaluate the performance of our algorithm against non-private ICL on standard benchmarks. Our experiments suggest that privacy protection is achievable with minimal loss in the performance. In some cases, such as for the TREC dataset [37] (c.f. Table 1), our algorithm can generate synthetic 4-shot demonstrations with a strong privacy level $\epsilon=1$ for ICL, achieving 50.7% accuracy with GPT-3 Babbage model. While substantially improving on the fully private 0-shot 35.4% accuracy, it is competitive even with the non-private solution, which achieves 50.6% accuracy.
  • While we focus on ICL with formal DP guarantees that use private data, we also explore approaches that do not make use of private demonstrations at all. We ask: can we improve on the zero-shot performance of models by asking the model to generate demonstrations required for ICL and then using it for solving the downstream task? Note that in this approach we do not use the private dataset at all. We present some interesting findings on this topic, which can be of independent interest in the non-private domain. We show that in some applications, LLMs can indeed generate relevant few-shot demonstrations on their own with pure instructions and perform well without needing any private data. For example, for AGNews dataset [44], GPT-3 Babbage model can generate 4-shot demonstrations solely with an instruction not using any data samples from the dataset and this achieves 68.0% ICL performance. This is competitive with the non-private solution, which achieves 69.3% accuracy. On the other hand, zero-shot performance of the model is only 47.9%. This suggests that privacy may come for free in certain tasks, when a model can generate its own demonstrations instead of answering the questions directly.
Independently and concurrently with our work, [30] and [14] also studied ICL with DP guarantees. While the motivations are similar, there are several differences between our work and theirs. We provide a detailed comparison in Section 6.
@@ 2.1 Prompting and In-context Learning
In-context learning (ICL) involves task adaptation without modifying a pre-trained model’s weights. Instead, ICL conditions the model by utilizing a series of examples (also called demonstrations). An example typically consists of an input and its corresponding ground-truth label output, both transformed into a specific format using a pattern and a verbalizer. Sequential demonstrations are then provided to the model, followed by a modified test input through pattern transformation. An instruction can also be added to provide the model with a set of output labels, or a description of the task. The model’s objective is to predict the label for this final data point. As LLMs gain more capability, ICL has emerged as a popular approach in natural language processing (NLP) tasks [12].
@@ 2.2 Differential Privacy
Definition 2.1 (Differential Privacy (DP) [16]).
@@
DP-SGD [1] and PATE [31] are the two main approaches of how DP can be integrated into model training. The DP-SGD algorithm modifies the standard SGD algorithm by per-sample clipping of gradients and adding carefully calibrated noise to the gradient updates during each iteration of training. On the other hand, PATE performs non-private training for a number of teacher models on disjoint subsets of the private dataset. The final student model is trained through privately aggregated predictions of teacher models to achieve DP guarantees.
@@ 3 Problem Definition, Threat Model, and Notations
Input to our problem are privacy parameters $\epsilon>0$, $\delta\in[0,1]$, a private dataset $\mathcal{D}_{\text{priv}}=\{z_{1},\cdots,z_{n}\}$ where each $z_{i}$ may consist of a content and a label part: $z_{i}=(x_{i},y_{i})$. We define two databases as neighboring if they differ by a single entry $z_{i}$. We are also given a pre-trained language model, $LM(x_{n}\mid x_{1},\cdots,x_{n-1})\in[0,1]$, where each $x_{i}$ is a token in a vocabulary $\mathcal{V}$: $x_{i}\in\mathcal{V}$. We formally define a single instance ICL as the following function computation. We are given a query input $q$ and a set of demonstrations $C\subseteq\mathcal{D}_{\text{priv}}$. The final predicted output (label) $y$ by the language model $LM$ for the query $q$ is
@@
We solve this problem by generating synthetic few-shot examples $C=\{\widetilde{z}_{1},\cdots,\widetilde{z}_{{n}_{\text{shots}}}\}$ from the underlying distribution of $\mathcal{D}_{\text{priv}}$ while satisfying $(\epsilon,\delta)$-DP on the private dataset $\mathcal{D}_{\text{priv}}$ without fine-tuning ${LM}$ (Figure 2). From the post-processing property of DP [15], this ensures that output in Eq. (1) would also satisfy $(\epsilon,\delta)$-DP with infinite number of queries.
@@
As an example scenario, for the basic sentiment analysis task, $PB$ can construct the following prompt:
@@
where the first sentence is the instruction, the label $y$ is positive and we provide a single demonstration from $\mathcal{D}^{y}$ in this example. Each task has an associated $PB$, which is provided in Table 5 and 6 of Appendix B.
@@ 4 Proposed Method
Our framework for ICL with DP guarantees consists of two steps (Figure 2):
@@
    Use the synthetic few-shot generations done in step 1 as demonstrations during the inference (doing ICL for queries).
@@
Figure 3: Illustration of step 1 (DP few-shot generation) in our framework (Figure 2). The example shows a synthetic demonstration generated token by token for the topic school with DP. The operations in Algorithm 1 for one step of generation (the token College) are depicted step by step.
@@ 4.1 Algorithm for DP Few-shot Generation
Algorithm 1 presents the pseudo code of the proposed algorithm for the first step and Figure 3 provides a demonstration with an example. For a given label $y$, we generate one token at a time starting from an empty list. At each token generation, we randomly sample $M\times N$ examples from the private dataset $\mathcal{D}_{\text{priv}}$ and partition them into $M$ disjoint subsets of $N$ samples each (lines 1-1). Each subset is appended with what is generated so far ($+$ is used for concatenation) and contributes with the probability of generating the next token (line 1). Before private aggregation, the effect of noise can be reduced by limiting the vocabulary to the tokens that are present in top-$K$ indices of the next-token probability coming from only the instruction without using any private data (lines 1-1). [^2] By sampling $K$, rather than $|V|$, random numbers, we significantly reduce the probability we sampled an outlier as $K\ll|V|$; the same technique was used in [34]. Next-token generation probabilities obtained from each subset are then privately aggregated. We consider both the Gaussian DP mechanism (line 1) and Report-Noisy-Max with Exponential DP mechanism (lines 1-1). Finally, the next token is produced and appended to what is generated so far. This continues until we either get the special end of sequence (EOS) token or hit the limit of maximum number of tokens we would like to generate.
@@ 4.2 Privacy Analysis
We defer a full proof of the theorem below to Appendix A, and only give high-level ideas of the analysis.
@@
Algorithm 1 is $(\epsilon,\delta)$ differentially private.
@@
Our proof follows the framework used in analyzing DP-SGD algorithm [1, 21], which consists of 3 steps. First, we show that each iteration of the algorithm is $(\epsilon,\delta)$-DP with some privacy parameters: For the Gaussian mechanism this follows from bounding the $\ell_{2}$-sensitivity of the term in line 1. For the variant using Exponential noise we appeal to the Report-Noisy-Max with Exponential Noise mechanism [15]. Next, we argue that as our algorithm draws $MN$ samples from a dataset of size $|\mathcal{D}_{\text{priv}}|$, there is a privacy amplification by subsampling at each iteration. Finally, we appeal to numerical composition theorems (Theorem A.3) to compose privacy loss across all the $T_{\max}$ iterations of our subsampled DP algorithms. ∎
@@ 5.1 Implementation Settings
In our experimentation, we follow the prior work on ICL [45] and use the following setup. [^3] Our repository is located at https://github.com/microsoft/dp-few-shot-generation.
Datasets
We study 4-way news classification AGNews [44], 6-way question classification TREC [37], and 14-way topic classification DBPedia [44] datasets for text classification tasks. For information extraction tasks, we study a slot filling dataset MIT Movies trivia10k13 [24], which consists of movie genre (MIT-G) and director name (MIT-D) as the slots to be filled. Further details about the datasets are provided in Appendix C.
The prompt construction functions used during synthetic few-shot generations in Algorithm 1 are provided with examples for each dataset in Appendix B. We use the synthetic few-shot generations $\{\widetilde{z}_{1},\cdots,\widetilde{z}_{{n}_{\text{shots}}}\}$ from Algorithm 1 as demonstrations in ICL for the aforementioned downstream tasks. We use the same prompt format during ICL following prior work [45] (see Tables 7-8 in Appendix D). We use the calibration approach in Zhao et al. [45] by default unless otherwise stated.
We present our main results on GPT-3 Babbage model using the OpenAI API. We generate and use 4-shot demonstrations (randomly without replacement from the label set) for ICL. Ablation studies on varying numbers of shots, model size, and hyperparameters are deferred to Section 5.3 and Appendix E.
@@ 5.2 Main Results
Table 1: 4-shot ICL performance on the test set of downstream tasks with various baselines. The mean and standard deviation of accuracy is reported over 5 runs with different random seeds. $\epsilon=0$ (0-shot) and $\epsilon=0$ (4-shot) are fully private solutions. The latter uses purely an instruction to generate synthetic few-shot demonstrations for ICL. Our private solution with various privacy levels $\epsilon=1,2,4,$ and $8$ uses Algorithm 1 to generate DP synthetic few-shot demonstrations for ICL. $\epsilon=\infty$ is the non-private solution that reports the maximum of two performances for ICL: (i) using generations of our Algorithm 1 with $\sigma=0$ and (ii) using randomly chosen 4-shot samples from the private data. Our private solution, in general, provides substantial gains over the fully private baselines and achieves competitive results with the non-private cases.
@@
We present our main results [^4] MIT-G and MIT-D results are uncalibrated accuracies as we did not observe an improvement with calibration. in Table 1 for various privacy levels. We fix $\delta=1/|\mathcal{D}_{\text{priv}}|$, however, with the advantage of privacy curves of DP mechanisms described in Appendix A smaller $\delta$ is guaranteed with minor increase in $\epsilon$ (e.g., Algorithm 1 can be both ($\epsilon=1,\delta=3\times 10^{-4}$)-DP and ($\epsilon=1.9,\delta=3\times 10^{-5}$)-DP for AGNews dataset). We provide the mean and the standard deviation of the accuracy on the test data with ICL over 5 runs with different random seeds.
@@
We observe that compared to 0-shot performance, which is a fully private solution, our approach provides substantial gains even with strict privacy levels as small as $\epsilon=1$. Additionally, our approach is competitive with the non-private solution ($\epsilon=\infty$), therefore, enables similar performance while providing strong privacy guarantees for the original data samples. Note that we consider two approaches for $\epsilon=\infty$ column; (i) running Algorithm 1 with $\sigma=0$ and (ii) 4-shot demonstrations randomly selected from the private dataset. We present best of two accuracies between the two approaches. Notably, (i) performs better for AGNews, DBPedia, TREC, and MIT-G whereas (ii) performs better for MIT-D.
@@
$\epsilon=0$ (4-shot) column presents an interesting finding which can be of independent interest in the non-private domain. To disambiguate the benefits of private data from LM’s own capability to generate relevant synthetic few-shot demonstrations, we run the following experiment. We generate synthetic 4-shot samples $\{\widetilde{z}_{1},\cdots,\widetilde{z}_{4}\}$ using purely the instruction without any private data (i.e. running only line 1 and auto-regressively generating the synthetic samples in Algorithm 1). Using these samples in ICL over the test data is reported in $\epsilon=0$ (4-shot) column. We point out that the model can substantially improve its 0-shot performance by generating relevant few-shot demonstrations with its existing capability. This can go to the extent that it can be fully competitive with even non-private performance for AGNews data. We believe that this behavior is consistent with the recent chain-of-thought prompting [39] and Bayesian inference framework for ICL [40, 28]. However, we do observe a significant improvement when $\{\widetilde{z}_{1},\cdots,\widetilde{z}_{4}\}$ are generated with the help of private data in Algorithm 1 for all the other datasets, which shows the benefits of using private data in ICL framework.
@@
Finally, we provide the parameters that we use for our main results in Table 9 and the hyperparameter search for all the datasets in Table 10-14 respectively in Appendix E. We state here a number of important observations. First of all, when we reduce the vocabulary to top-100 tokens using public instruction during generation (i.e. RVP is True with $K=100$ in Algorithm 1), we do not observe high variation among the choices of $M,N$. This is good in terms of the stability of our algorithm. We notice that showing $N>1$ examples per subset helps in general although the best signal-to-noise ratio is achieved when $N=1$ for a fixed $MN$. This is possibly due to reducing the effect of noise by limiting top-100 tokens from the public instruction. When we set RVP as False in Algorithm 1, we add noise to the full vocabulary during generation, which results in higher variation among the choices of $M,N$. As expected, the best performance is in general achieved for highest $MN$ and $N=1$ as this setting is the most optimal for signal-to-noise ratio. Finally, we observe that setting RVP as False matches the performance of setting RVP as True in Algorithm 1 with largest $MN$ but that also comes with more queries to the LM per token generation.

### v2

@@ 1 Introduction
The emergence of in-context learning (ICL) with large language models (LLMs), popularized by the seminal work of Brown et al. (2020), has revolutionized the field of natural language processing and machine learning; see Dong et al. (2023) for a survey on ICL and the references therein. In-context learning involves downstream task adaptation without modifying a pre-trained model’s weights. This is achieved by conditioning the model through a series of demonstrations of the task at hand appended as a prompt. An advantage of ICL is that it offers a cost-effective and adaptable alternative to fine-tuning LLMs. By leveraging the model’s pre-trained knowledge, it enables efficient generalization across tasks, allows for quick adaptation to new domains or concepts, and requires only a handful of labeled examples for adaptation.
However, privacy is a concern when deploying LLMs with users’ data incorporated into prompts. As an example, consider healthcare AI applications, where clinical reports belonging to the patients may be used as demonstrations to provide relevant context to the LLM to answer queries. A malicious adversary might attempt to circumvent API restrictions through jailbreaking thereby gaining direct access to the demonstrations as depicted in Fig. 1.111As also observed in a real-world system built on an LLM: see https://tinyurl.com/nhzadhuz. More generally, it is a major concern that LLMs may regurgitate prompt data in their output (Priyanshu et al., 2023; Duan et al., 2023; Wang et al., 2023). These scenarios raise privacy risks regarding the data used for constructing the prompt. To mitigate these risks associated with prompt data, heuristic precautions can be taken such as removing personal identifiable information (PII). Unfortunately, it is well known in the privacy literature that a document without any PII may still be linked to an individual with auxiliary information an adversary possesses, which will inevitably lead to privacy violations (e.g., based on GDPR (Art. 29 WP, 2014)).
In this context, differential privacy (DP) (Dwork et al., 2006b) is widely regarded as the gold standard in safeguarding the privacy of individuals contributing their data to various applications, ensuring that their sensitive information remains confidential while still enabling valuable insights to be derived from the aggregated data (US Census Bureau, 2020).
Figure 1: Description of a potential privacy violation when few-shot demonstrations are pulled from a private dataset in an ICL framework for a healthcare application. A malicious adversary attempts a basic prompt injection attack and gains direct access to the demonstrations. Basic heuristics such as personal identifiable information (PII) removal may still leave linkable information (Art. 29 WP, 2014) to an individual in case the adversary has auxilary information (e.g., a unique patient with a particular disease or treatment) and do not prevent against privacy violations.
@@ 1.1 Our Contributions
  • We introduce a new algorithm to perform ICL with DP guarantees on private datasets (Fig. 2). Our algorithm generates synthetic few-shot demonstrations from the original private dataset to be used for ICL during inference, and guarantees DP with respect to examples in the private dataset. At a high level, our approach utilizes the capabilities of LLMs in terms of generating a data sample similar to the ones in the original dataset based on demonstrations of a number of samples from the private data. To guarantee DP on the private dataset, we perform this operation by privately aggregating the generation process coming across disjoint subsets of the private data.
  • We empirically evaluate the performance of our algorithm against non-private ICL on standard benchmarks. Our experiments suggest that privacy protection is achievable with minimal loss in the performance. In some cases, such as for TREC dataset (c.f. Tab. 1), our algorithm can generate synthetic 4-shot demonstrations with a strong privacy level $\epsilon=1$ for ICL, achieving 50.7% accuracy with GPT-3 Babbage model. While substantially improving on the fully private 0-shot 35.4% accuracy, it is competitive even with the non-private solution, which achieves 50.6% accuracy.
  • While we focus on ICL with formal DP guarantees that use private data, we also explore approaches that do not make use of private demonstrations at all. We ask: can we improve on the zero-shot performance of models by asking the model to generate demonstrations required for ICL and then using it for solving the downstream task? Note that in this approach we do not use the private dataset at all. We present some interesting findings on this topic, which can be of independent interest in the non-private domain. We show that in some applications, LLMs can indeed generate relevant few-shot demonstrations on their own with pure instructions and perform well without needing any private data. For example, for AGNews dataset (Zhang et al., 2015), GPT-3 Babbage model can generate 4-shot demonstrations solely with an instruction not using any data samples from the dataset and this achieves 68.0% ICL performance. This is competitive with the non-private solution, which achieves 69.3% accuracy. On the other hand, zero-shot performance of the model is only 47.9%. This suggests that privacy may come for free in certain tasks, when a model can generate its own demonstrations instead of answering the questions directly.
Independently and concurrently with our work, Wu et al. (2024) and Duan et al. (2023) also studied ICL with DP guarantees. While the motivations are similar, there are several differences between our work and theirs. We provide a detailed comparison in Section 6.
@@ 2.1 Prompting and In-context Learning
In-context learning (ICL) involves task adaptation without modifying a pre-trained model’s weights. Instead, ICL conditions the model by utilizing a series of examples (also called demonstrations). An example typically consists of an input and its corresponding ground-truth label, both transformed into a specific format using a pattern and a verbalizer. Sequential demonstrations are then provided to the model, followed by a modified test input through pattern transformation. An instruction can also be added to provide the model with a set of output labels, or a description of the task. The model’s objective is to predict the label for this final data point. As LLMs gain more capability, ICL has emerged as a popular approach in natural language processing (NLP) tasks (Dong et al., 2023).
@@ 2.2 Differential Privacy
Definition 2.1 (Differential Privacy (DP) (Dwork et al., 2006a)).
@@
DP-SGD (Abadi et al., 2016) and PATE (Papernot et al., 2017) are the two main approaches of how DP can be integrated into model training. The DP-SGD algorithm modifies the standard SGD algorithm by per-sample clipping of gradients and adding carefully calibrated noise to the gradient updates during each iteration of training. On the other hand, PATE performs non-private training for a number of teacher models on disjoint subsets of the private dataset. The final student model is trained through privately aggregated predictions of teacher models to achieve DP guarantees.
@@ 3 Problem Definition, Threat Model, and Notations
Input to our problem are privacy parameters $\epsilon>0$, $\delta\in[0,1]$, a private dataset $\mathcal{D}_{\text{priv}}=\{z_{1},\cdots,z_{n}\}$ where each $z_{i}$ may consist of a content and a label part: $z_{i}=(x_{i},y_{i})$. We define two databases as neighboring if they differ by a single entry $z_{i}$. We are also given a pre-trained language model, $LM(x_{n}\mid x_{1},\cdots,x_{n-1})\in[0,1]$, where each $x_{i}$ is a token in a vocabulary $\mathcal{V}$: $x_{i}\in\mathcal{V}$. We formally define a single instance ICL as the following function computation. We are given a query input $q$ and a set of demonstrations $C\subseteq\mathcal{D}_{\text{priv}}$. The final predicted output (label) $y$ by the $LM$ for the query $q$ is
@@
We solve this problem by generating synthetic few-shot examples $C=\{\widetilde{z}_{1},\cdots,\widetilde{z}_{{n}_{\text{shots}}}\}$ from the underlying distribution of $\mathcal{D}_{\text{priv}}$ while satisfying $(\epsilon,\delta)$-DP on the private dataset $\mathcal{D}_{\text{priv}}$ without fine-tuning ${LM}$ (Fig. 2). From the post processing property of DP (Dwork & Roth, 2014), this ensures that output in Eq. (1) would also satisfy $(\epsilon,\delta)$-DP with infinite number of queries.
@@
As an example scenario, for basic sentiment analysis task, $PB$ can construct the following prompt:
@@
where the first sentence is the instruction, the label $y$ is positive and we provide a single demonstration from $\mathcal{D}^{y}$ in this example. Each task has an associated $PB$, provided in Tab.5 and 6 of Appendix B.
@@ 4 Proposed Method
Our framework for ICL with DP guarantees consists of two steps (Fig. 2):
@@
    Use the synthetic few-shot generations from step 1 as ICL demonstrations during the inference.
@@
Figure 3: Illustration of step 1 (DP few-shot generation) in our framework (Fig. 2). The example shows a synthetic demonstration generated token by token for the topic school with DP. The operations in Alg. 1 for one step of generation (the token College) are depicted step by step.
@@ 4.1 Algorithm for DP Few-shot Generation
Alg. 1 presents the pseudo code of the proposed algorithm for the first step and Fig. 3 provides a demonstration with an example. For a given label $y$, we generate one token at a time starting from an empty list. At each token generation, we randomly sample $MN$ examples from the private dataset $\mathcal{D}_{\text{priv}}$ and partition them into $M$ disjoint subsets of $N$ samples each (lines 1-1). Each subset is appended with what is generated so far ($+$ is used for concatenation) and contributes with the probability of generating the next token (line 1). Before private aggregation, the effect of noise can be reduced by limiting the vocabulary to the tokens that are present in top-$K$ indices of the next-token probability coming from only the instruction without using any private data (lines 1-1). [^2] By sampling $K$, rather than $|V|$, random numbers, we significantly reduce the probability we sampled an outlier as $K\ll|V|$; the same technique was used in Tian et al. (2022). Next-token generation probabilities obtained from each subset are then privately aggregated. We consider both the Gaussian mechanism (line 1) and Report-Noisy-Max with Exponential mechanism (lines 1-1). Finally, the next token is produced and appended to what is generated so far. This continues until we either get the special end of sequence (EOS) token or hit the limit of maximum number of tokens we would like to generate.
@@ 4.2 Privacy Analysis
We defer a full proof of the theorem below to Appendix A, and give high-level ideas of the analysis.
@@
Alg. 1 is $(\epsilon,\delta)$ differentially private.
@@
Our proof follows the framework used in analyzing DP-SGD algorithm (Abadi et al., 2016; Gopi et al., 2021), which consists of 3 steps. First, we show that each iteration of the algorithm is $(\epsilon,\delta)$-DP with some privacy parameters: For the Gaussian mechanism this follows from bounding the $\ell_{2}$-sensitivity of the term in line 1. For the variant using Exponential noise we appeal to the Report-Noisy-Max with Exponential Noise mechanism (Dwork & Roth, 2014). Next, we argue that as our algorithm draws $MN$ samples from a dataset of size $|\mathcal{D}_{\text{priv}}|$, there is a privacy amplification by subsampling at each iteration. Finally, we appeal to numerical composition theorems (Theorem A.3) to compose privacy loss across $T_{\max}$ iterations of our subsampled DP algorithms. ∎
@@ 5.1 Implementation Settings
We follow the prior ICL work (Zhao et al., 2021) and use the following setup. [^3] Our repository is located at https://github.com/microsoft/dp-few-shot-generation.
Datasets. We study 4-way news classification AGNews (Zhang et al., 2015), 6-way question classification TREC (Voorhees & Tice, 2000), and 14-way topic classification DBPedia (Zhang et al., 2015) datasets for classification tasks. For information extraction tasks, we study a slot filling dataset MIT Movies trivia10k13 (Liu et al., 2012), which consists of movie genre (MIT-G) and director name (MIT-D) as the slots to be filled. Further details about the datasets are provided in Appendix C.
The prompt construction functions used in synthetic few-shot generations of Alg. 1 are provided with examples for each dataset in Appendix B. We use the synthetic few-shot generations $\{\widetilde{z}_{1},\cdots,\widetilde{z}_{{n}_{\text{shots}}}\}$ from Alg. 1 as demonstrations in ICL for the downstream tasks above. We use the prompt format during ICL following prior work (Zhao et al., 2021) (see Tab. 7-8 in Appendix D). We use the calibration approach in Zhao et al. (2021) by default unless otherwise stated.
We present our main results on GPT-3 Babbage using the OpenAI API. We generate and use 4-shot demonstrations (randomly w/out replacement from the label set) for ICL. Ablation studies on varying numbers of shots, model size, and hyperparameters are deferred to Section 5.3 and Appendix E.
@@ 5.2 Main Results
We present our main results [^4] MIT-G and MIT-D results are uncalibrated accuracies as we did not observe an improvement with calibration. in Tab. 1 for various privacy levels. We fix $\delta=1/|\mathcal{D}_{\text{priv}}|$, however, with the advantage of privacy curves of DP mechanisms described in Appendix A smaller $\delta$ is guaranteed with minor increase in $\epsilon$ (e.g., Alg. 1 can be both ($\epsilon=1,\delta=3\times 10^{-4}$)-DP and ($\epsilon=1.9,\delta=3\times 10^{-5}$)-DP for AGNews dataset). We provide the mean and the standard deviation of the accuracy on the test data with ICL over 5 runs with different random seeds.
Table 1: 4-shot ICL performance on the test set of downstream tasks with various baselines. $\epsilon=0$ (0-shot) and $\epsilon=0$ (4-shot) are fully private solutions. The latter uses purely an instruction to generate synthetic few-shot demonstrations for ICL. Our private solution with various privacy levels $\epsilon=1,2,4,$ and $8$ uses Alg. 1 to generate DP synthetic few-shot demonstrations for ICL. $\epsilon=\infty$ is the non-private solution that reports the maximum of two performances for ICL: (i) using generations of our Alg. 1 with $\sigma=0$ and (ii) using randomly chosen 4-shot samples from the private data. Our private solution, in general, provides substantial gains over the fully private baselines and achieves competitive results with the non-private cases.
@@
We observe that compared to 0-shot performance, which is a fully private solution, our approach provides substantial gains even with strict privacy levels as small as $\epsilon=1$. Additionally, our approach is competitive with the non-private solution ($\epsilon=\infty$), therefore, enables similar performance while providing strong privacy guarantees for the original data samples. Note that we consider two approaches for $\epsilon=\infty$ column; (i) running Alg. 1 with $\sigma=0$ and (ii) 4-shot demonstrations randomly selected from the private dataset. We present best of two accuracies between the two approaches. Notably, (i) performs better than (ii) for all the datasets except MIT-D.
@@
$\epsilon=0$ (4-shot) column presents an interesting finding which can be of independent interest in the non-private domain. To disambiguate the benefits of private data from LM’s own capability to generate relevant synthetic few-shot demonstrations, we run the following experiment. We generate synthetic 4-shot samples $\{\widetilde{z}_{1},\cdots,\widetilde{z}_{4}\}$ using purely the instruction without any private data (i.e. running only line 1 and auto-regressively generating the synthetic samples in Alg. 1). Using these samples in ICL over the test data is reported in $\epsilon=0$ (4-shot) column. We point out that the model can substantially improve its 0-shot performance by generating relevant few-shot demonstrations with its existing capability. This can go to the extent that it can be fully competitive with even non-private performance for AGNews data. We believe that this behavior is consistent with the recent chain-of-thought prompting (Wei et al., 2022) and Bayesian inference framework for ICL (Xie et al., 2022; Min et al., 2022). However, we do observe a significant improvement when $\{\widetilde{z}_{1},\cdots,\widetilde{z}_{4}\}$ are generated with the help of private data in Alg. 1 for all the other datasets, which shows the benefits of using private data in ICL framework.
@@
We provide the parameters that we use for our main results in Tab. 9 and the hyperparameter search in Tab. 10-14 respectively in Appendix E. We state here a number of important observations. First of all, when we reduce the vocabulary to top-100 tokens using public instruction during generation (i.e. RVP is True with $K=100$ in Alg. 1), we do not observe high variation among the choices of $M,N$. This is good in terms of the stability of Alg. 1. We notice that showing $N>1$ examples per subset helps in general although the best signal-to-noise ratio is achieved when $N=1$ for a fixed $MN$. This is possibly due to reducing the effect of noise by limiting top-100 tokens from the public instruction. When we set RVP as False in Alg. 1, we add noise to the full vocabulary during generation, which results in higher variation among the choices of $M,N$. As expected, the best performance is in general achieved for highest $MN$ and $N=1$ as this setting is the most optimal for signal-to-noise ratio. Finally, we observe that setting RVP as False matches the performance of setting RVP as True in Alg. 1 with largest $MN$ but that also comes with more queries to the LM per token generation.